### v2

# peapods: A Rust-Accelerated Monte Carlo Package for Ising Spin Systems

Abstract
We present peapods [1] (github.com/PeaBrane/peapods), an open-source Python package for Monte Carlo simulation of Ising spin systems with arbitrary coupling constants on periodic Bravais lattices with user-specified neighbor offsets. The computational core is written in Rust and exposed to Python via PyO3 [2], combining the ergonomic interface of Python with the performance of compiled, memory-safe code. The package implements Metropolis [3] and Gibbs single-spin-flip algorithms, Swendsen–Wang [4] and Wolff [5] cluster updates, parallel tempering [6], and three replica cluster moves for spin glasses: the Houdayer isoenergetic cluster move [7], the Jörg stochastic variant [8], and the Chayes–Machta–Redner (CMR) blue-bond algorithm [9]. Overlap statistics between replica pairs enable computation of the spin glass order parameter and Binder ratio. Replica-level parallelism is achieved through the Rayon [10] work-stealing scheduler. We validate the implementation against the exact critical temperatures of the two-dimensional Ising model on the square [11] and triangular [12, 13] lattices via finite-size scaling of the Binder cumulant [14].
1 Introduction
The Ising model is a cornerstone of statistical mechanics and the study of phase transitions. Despite its simplicity—binary spins on a lattice coupled by nearest-neighbor interactions—it exhibits rich critical phenomena and serves as a testbed for Monte Carlo algorithms. In the disordered-coupling case, it becomes the Edwards–Anderson spin glass [15], one of the central problems in condensed matter physics.
Monte Carlo simulations of spin systems are conceptually straightforward but computationally demanding. Near the critical temperature, critical slowing down causes single-spin-flip algorithms to decorrelate slowly, requiring cluster algorithms [4, 5] for efficient sampling. Disordered systems additionally benefit from parallel tempering [6] to escape metastable states, and from replica cluster moves such as the Houdayer isoenergetic cluster move [7] and its stochastic variants [8, 9] that exploit correlations between replicas at the same temperature to accelerate decorrelation in spin glasses.
Python is the dominant language for scientific computing, but its interpreter overhead makes it poorly suited for the tight inner loops of Monte Carlo simulation. Common mitigations include NumPy vectorization, Numba JIT compilation, or C/Fortran extensions. Each has drawbacks: vectorization requires materializing large temporary arrays, Numba has limited support for complex data structures, and C extensions sacrifice memory safety and developer productivity.
Rust offers an alternative: compiled performance competitive with C, strong memory safety guarantees enforced at compile time, and ergonomic tooling. The PyO3 library [2] and Maturin build tool [16] provide a mature ecosystem for writing Python extensions in Rust. In this work, we describe peapods, a package that places the entire sampling loop in Rust while exposing a clean Python interface for model construction, analysis, and visualization. The lattice geometry is fully general: users specify neighbor offsets as integer displacement vectors, enabling simulation on hypercubic, triangular, FCC, and other Bravais lattices.
2 Model
We consider the Ising Hamiltonian on a periodic Bravais lattice $\Lambda$ with $n$ neighbor directions:
$H=-\sum_{\langle i,j\rangle}J_{ij}\,\sigma_{i}\sigma_{j},$ (1)
where $\sigma_{i}\in\{-1,+1\}$ are the spin variables, the sum runs over nearest-neighbor pairs, and $J_{ij}$ are the coupling constants. The lattice geometry is defined by a shape (e.g. $L\times L$) and a set of $n$ integer offset vectors specifying the forward neighbor directions; hypercubic lattices use unit vectors along each axis ($n=d$), while the triangular lattice uses offsets $\{(1,0),(0,1),(1,-1)\}$ ($n=3$, coordination $z=6$). The couplings can be specified as an arbitrary user-provided array of shape $(\text{lattice\_shape},n)$, or via one of three built-in distributions:
  • Ferromagnetic: $J_{ij}=1$ for all bonds.
  • Bimodal: $J_{ij}\in\{-1,+1\}$ with equal probability (the $\pm J$ spin glass).
  • Gaussian: $J_{ij}\sim\mathcal{N}(0,1)$.
The package simulates $R$ independent replicas, each consisting of a full parallel tempering ladder across temperatures $T_{1}<T_{2}<\cdots<T_{K}$, giving $R\times K$ total systems. Multiple replicas at the same temperature enable replica cluster moves and the computation of overlap statistics for spin glass order parameters.
3 Algorithms
3.1 Single-spin-flip updates
Metropolis.
The Metropolis algorithm [3] proposes flipping each spin $\sigma_{i}\to-\sigma_{i}$ and accepts with probability
$P_{\text{accept}}=\min\!\big(1,\,e^{-\Delta E/T}\big),$ (2)
where $\Delta E=2\sigma_{i}h_{i}$ is the energy change and $h_{i}=\sum_{j\in\partial i}J_{ij}\sigma_{j}$ is the local field. The implementation avoids computing the exponential explicitly by comparing $\Delta E$ against $(T/2)\ln u$ where $u\sim\text{Uniform}(0,1)$.
Gibbs.
The Gibbs (heat bath) algorithm sets $\sigma_{i}=+1$ with probability $(1+e^{-2h_{i}/T})^{-1}$. This is implemented by comparing $\Delta E$ against the logit transform $(T/2)\ln[u/(1-u)]$.
Both sweeps visit every spin once in sequential order. The local field $h_{i}$ is computed on-the-fly from both forward and backward neighbors in each lattice dimension.
3.2 Cluster updates
Cluster algorithms exploit the Fortuin–Kasteleyn representation [17] to build correlated clusters of spins and flip them collectively, dramatically reducing autocorrelation times near $T_{c}$.
Swendsen–Wang.
The Swendsen–Wang algorithm [4] activates each satisfied bond (where $J_{ij}\sigma_{i}\sigma_{j}>0$) with probability $p=1-e^{-2J_{ij}\sigma_{i}\sigma_{j}/T}$. Activated bonds define clusters via a union-find data structure with path compression and union by rank. Each cluster is then flipped with probability $1/2$.
Wolff.
The Wolff algorithm [5] grows a single cluster from a random seed spin via breadth-first search. Each aligned neighbor is added with probability $p=1-e^{-2J_{ij}\sigma_{i}\sigma_{j}/T}$ when $J_{ij}\sigma_{i}\sigma_{j}>0$. The entire cluster is then flipped deterministically. While each Wolff step flips fewer spins than Swendsen–Wang, it avoids the overhead of a full union-find pass over the lattice.
3.3 Parallel tempering
Parallel tempering [6] improves ergodicity by allowing replicas at different temperatures to exchange configurations. At each tempering step, a random adjacent pair $(T_{k},T_{k+1})$ is selected and a swap is accepted with probability
$P_{\text{swap}}=\min\!\big(1,\,e^{\Delta}\big),\qquad\Delta=N(E_{k+1}-E_{k})\!\left(\frac{1}{T_{k}}-\frac{1}{T_{k+1}}\right),$ (3)
where $E_{k}$ is the average energy per spin of the replica currently at temperature $T_{k}$ and $N$ is the total number of spins. The swap is implemented by exchanging entries in a system_ids permutation array that maps temperatures to physical replica indices, avoiding any copying of spin data.
3.4 Replica cluster moves
The package implements three replica cluster algorithms for spin glasses, all operating on pairs of replicas at the same temperature. The common framework is:
  1. 1.
    For each temperature, randomly pair the replicas. For each pair $(a,b)$, identify an eligible subgraph based on the overlap $\sigma_{i}^{(a)}\sigma_{i}^{(b)}$.
  2. 2.
    Build clusters on this subgraph (via BFS for Wolff-style single-cluster, or union-find for Swendsen–Wang-style global decomposition).
  3. 3.
    Exchange the spins in selected clusters between the two replicas: $\sigma_{i}^{(a)}\leftrightarrow\sigma_{i}^{(b)}$ for all $i$ in the cluster.
Houdayer ICM.
The Houdayer isoenergetic cluster move [7] restricts to the negative-overlap sites $\mathcal{N}=\{i:\sigma_{i}^{(a)}\neq\sigma_{i}^{(b)}\}$ and activates all bonds deterministically (probability 1). This exchange is isoenergetic: because every bond crossing the cluster boundary connects a site inside $\mathcal{N}$ (where the replicas disagree) to a site outside $\mathcal{N}$ (where they agree), the local field at each boundary site is unchanged by the swap. The move is always accepted with no Metropolis test, and the cluster growth is temperature-independent, which is particularly valuable at low temperatures where single-spin-flip moves freeze out.
Jörg move.
Jörg [8] proposed a refinement for three-dimensional systems where Houdayer clusters can span the lattice. The eligible subgraph is the same (negative-overlap sites), but bonds are activated stochastically with probability $p=1-e^{-4J_{ij}\sigma_{i}^{(a)}\sigma_{j}^{(a)}/T}$ on satisfied bonds ($J_{ij}\sigma_{i}^{(a)}\sigma_{j}^{(a)}>0$), breaking up system-spanning clusters into smaller pieces.
CMR move.
The Chayes–Machta–Redner (CMR) algorithm [9] generalizes the overlap cluster approach by bonding all same-sign overlap sites (not just negative-overlap), with stochastic bond activation at probability $p=1-e^{-4J_{ij}\sigma_{i}^{(a)}\sigma_{j}^{(a)}/T}$ on satisfied bonds. This provides a complementary decorrelation channel to the Houdayer and Jörg moves.
All three modes support both Wolff (single-cluster BFS) and Swendsen–Wang (global union-find) scan styles, selectable via the overlap_cluster_mode parameter.
3.5 Overlap statistics
For spin glass systems, the relevant order parameter is the overlap between replica pairs:
$q=\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}^{(a)}\sigma_{i}^{(b)}.$ (4)
The package computes $\langle q\rangle$, $\langle q^{2}\rangle$, and $\langle q^{4}\rangle$ over the measurement sweeps for each consecutive replica pair at each temperature. From these, the spin glass Binder ratio [14]
$g=1-\frac{\langle q^{4}\rangle}{3\langle q^{2}\rangle^{2}}$ (5)
can be used for finite-size scaling analysis of the spin glass transition, analogously to the ferromagnetic Binder cumulant. The crossing of $g(T)$ for different system sizes locates the critical temperature $T_{c}$ [18].
4 Implementation
4.1 Architecture
The package follows a mixed Rust/Python layout built with Maturin [16]. The Rust library is compiled to a shared object (peapods._core) and imported by a thin Python layer that handles model construction, post-processing of observables (Binder cumulant, heat capacity), and visualization.
The central Rust struct is IsingSimulation, exposed to Python as a PyO3 pyclass. It owns:
  • A flat Vec<i8> of all spin configurations across replicas ($R\times N$ entries).
  • A flat Vec<f32> of coupling constants ($N\times d$ entries, storing only forward couplings—backward couplings are derived on-the-fly as the forward coupling of the backward neighbor).
  • Per-replica Xoshiro256** random number generators for thread-safe parallel execution.
  • A system_ids permutation vector for parallel tempering bookkeeping.
4.2 Lattice geometry
The Lattice struct supports arbitrary Bravais lattice geometries specified by a shape (extents along each axis) and a set of $n$ integer offset vectors defining the forward neighbor directions. For hypercubic lattices, these are the $d$ unit vectors; for the triangular lattice in 2D, they are $\{(1,0),(0,1),(1,-1)\}$. At construction time, the full neighbor table is precomputed: for each of the $N$ sites and each of the $n$ directions, both the forward and backward neighbor flat indices are stored, giving a table of $2nN$ entries (u32). Periodic boundary conditions are applied via rem_euclid on the lattice coordinates.
This precomputed table replaces per-call division and modular arithmetic with a single array lookup, which is faster in practice. The memory cost is modest: for example, 96 KB for a $64\times 64$ triangular lattice ($n=3$) or 6 MB for a $64^{3}$ cubic lattice ($n=3$).
4.3 Parallelism
All replicas are independent conditioned on the temperature assignment, so single-spin-flip sweeps, energy computations, and cluster updates are trivially parallelizable across replicas. We use Rayon’s [10] work-stealing thread pool via a shared par_over_replicas helper that dispatches a closure over replicas in parallel. Each replica receives a disjoint mutable slice of the spin array and its own RNG, avoiding synchronization.
The one exception is parallel tempering, which is inherently sequential (swaps involve pairs of replicas sharing temperature-space neighbors). This step is cheap relative to the sweeps and does not limit scalability.
4.4 Memory management
All buffers (spins, couplings, energies, interactions) are allocated once at initialization and reused across sweeps. The only per-sweep allocations are within cluster updates: union-find arrays and Wolff’s BFS stack, which are allocated per-replica inside the parallel closure. Running statistics are accumulated in f64 to avoid precision loss over many sweeps.
5 Validation
We validate the implementation using exactly solvable two-dimensional ferromagnetic Ising models.
The Binder cumulant [14],
$U_{L}=1-\frac{\langle m^{4}\rangle}{3\langle m^{2}\rangle^{2}},$ (6)
where $m=N^{-1}\sum_{i}\sigma_{i}$ is the magnetization per spin, is a dimensionless ratio that crosses at $T_{c}$ for different system sizes $L$, with the crossing value determined by the universality class.
Square lattice.
Figure 1 shows $U_{L}(T)$ for $L=8,16,32,64$ on the square lattice computed with 50,000 sweeps of Metropolis updates combined with Swendsen–Wang cluster updates and parallel tempering. All four curves cross at $T\approx 2.27$ with $U\approx 0.61$, consistent with the exact critical temperature $T_{c}=2/\ln(1+\sqrt{2})\approx 2.269$ [11] and the universal Binder cumulant value of the 2D Ising universality class.
Triangular lattice.
As a test of the general Bravais lattice geometry, we repeat the analysis on the triangular lattice with neighbor offsets $\{(1,0),(0,1),(1,-1)\}$ (coordination $z=6$). The exact critical temperature is $T_{c}=4/\ln 3\approx 3.641$ [12, 13]. For $L=8,16,32$ with 5,000 sweeps, the Binder cumulants cross at $T\approx 3.64$ with $U\approx 0.61$, confirming both the correctness of the neighbor table and the universality of the crossing value (same 2D Ising universality class).
Figure 1: Binder cumulant $U_{L}$ vs. temperature for $L=8,16,32,64$ on the square lattice. The dashed line marks the exact critical temperature $T_{c}\approx 2.269$. All four sizes cross near $U\approx 0.61$.
6 Performance
By moving the entire sampling loop to compiled Rust code, the package avoids the overhead inherent in Python-level Monte Carlo implementations, including per-sweep temporary array allocations, interpreter dispatch in tight loops, and the cost of Python-level iteration over replicas for cluster updates. The Rust backend additionally benefits from on-the-fly neighbor computation (avoiding memory indirection through neighbor tables), Rayon work-stealing parallelism across replicas, and the Xoshiro256** generator which is substantially faster than NumPy’s default RNG for scalar draws. Together, these yield significant speedups over equivalent Python/NumPy implementations, particularly for cluster algorithms where per-temperature Python loops are the bottleneck.
7 Usage
The Python interface is designed for simplicity:
import numpy as np
from peapods import Ising
# Square lattice ferromagnet
model = Ising((32, 32), couplings="ferro",
temperatures=np.geomspace(0.1, 10, 32))
model.sample(10000, sweep_mode="metropolis",
cluster_update_interval=1, cluster_mode="sw",
pt_interval=1)
print(model.binder_cumulant)
# Triangular lattice ferromagnet
tri = Ising((32, 32), temperatures=np.linspace(3.0, 4.2, 32),
n_replicas=2,
neighbor_offsets=[[1, 0], [0, 1], [1, -1]])
tri.sample(5000, sweep_mode="metropolis",
cluster_update_interval=1, pt_interval=1)
print(tri.binder_cumulant)
# Spin glass with Houdayer ICM
sg = Ising((8, 8, 8), couplings="bimodal",
temperatures=np.linspace(0.8, 1.4, 24),
n_replicas=4)
sg.sample(10000, sweep_mode="metropolis",
pt_interval=1, houdayer_interval=1)
print(sg.sg_binder)
The package supports arbitrary Bravais lattice geometries via the neighbor_offsets parameter (defaulting to hypercubic when omitted) and all three coupling types. Installation from PyPI is via pip install peapods, with pre-built wheels for Linux (x86_64, aarch64), macOS (Intel, Apple Silicon), and Windows (x86_64).
8 Conclusion
We have presented peapods, a Monte Carlo simulation package for Ising spin systems that combines a Python interface with a Rust computational backend. The package implements the standard toolkit of Metropolis, Gibbs, Swendsen–Wang, Wolff, and parallel tempering algorithms, three replica cluster moves for spin glasses (Houdayer ICM, Jörg, and CMR), overlap statistics for spin glass order parameters, and support for arbitrary Bravais lattice geometries via user-specified neighbor offsets. Replica-level parallelism is achieved via Rayon. The implementation is validated against exact 2D Ising critical points on both the square and triangular lattices. The package is open-source and available on PyPI.
Several directions for future work are planned:
  • Generalized spin models: extending beyond Ising ($\mathbb{Z}_{2}$) spins to support $q$-state Potts, $\mathbb{Z}_{q}$ clock, and continuous O($N$) models, including their disordered variants.
  • Cluster algorithms for frustrated systems: implementing algorithms such as the KBD algorithm that extend cluster updates to systems with competing interactions where standard Swendsen–Wang and Wolff methods are inapplicable.